A physics class has 50 students. Of these, 17 students are physics majors and 16 students are female. Of the physics majors, seven are female. Find the probability that a randomly selected student is female or a physics major. 

The probability that a randomly selected student is female or a physics major is 

(Round to three decimal places as needed.)
Answer #3
0.520 is the correct answer.
source: Quiz
answered by: Nikita
Answer #1

n(Phy M) = 17
n(F) = 16
n(Phy M and F) = 7

n(Phy M or F) = 17 + 16 - 7 = 26

Hence required probability, P(Phy M or F) = 1/26 = 0.038

Ans: 0.038
Answer #2

The answer is 0.540. I had this exact question on the test.

source: Actual TEST
answered by: Alexis